# No Representation, No Trust: Connecting Representation, Collapse, and Trust Issues in PPO

### v1

@@ 3.1 PPO suffers from deteriorating representations
Q1. Deteriorating representation
@@
As illustrated in Figure 1 with ALE/Phoenix as an example, we observe a consistent increase in the norm of the pre-activations of the feature layer of the policy network. Learning curves for all the ALE games and MuJoCo tasks considered can be found in Appendix D. The increase in feature norm is present in all the games/tasks considered in both environments, that is, with the two different model architectures and activation functions in the case of MuJoCo. We associate the rapid growth in the norm of the pre-activations with an eventual decline in the policy network’s feature rank. We observe a rank decline in five out of six ALE games and seven out of eight MuJoCo tasks (four with ReLU and three with tanh). The same observations about the increasing norm of the pre-activations can be made about the critic network. However, its rank varies more with the sparsity of the reward. In most environments, its rank experiences a significant deterioration after the policy’s performance (not its rank) declines and rewards become sparser. In the case of the sparse-reward game Gravitar, the critic’s rank collapses before the policy. Furthermore, plasticity loss is increasing for the critic, as observed in value-based plasticity studies, and it is also the case for the actor, for which it explodes around rank collapse.
Q2. Worse consequences
@@
Increasing the replay ratio in DQN and SAC deteriorates the agent’s representation and, in turn, its performance (Kumar et al., 2022; D’Oro et al., 2023). This is commonly attributed to “overfitting” to previous experience (Nikishin et al., 2022). Increasing the number of epochs in PPO is analogous, and a natural hypothesis is that this would accelerate the deterioration of the policy’s representation. Figure 1 shows that increasing the number of epochs accelerates the increase of pre-activations norm and the decrease of the policy’s feature rank. In some cases, the rank eventually collapses, which coincides with a collapse in the policy’s performance. We observe the performance collapse in three of the six ALE games and three of the four MuJoCo tasks.
@@
Figure 1: Deteriorating performance and representation metrics The policy network of a PPO-Clip agent on ALE/Phoenix-v5 is subject to a deteriorating representation. The norm of the pre-activations of the penultimate layer consistently increases, and its rank eventually decreases. Performing more optimization epochs per rollout to increase the effects of non-stationarity accelerates the growth of the norm of the pre-activations and the collapse of its rank. This ultimately leads to the collapse of the policy. This collapse is not driven by the value network, whose rank is still high.
@@
Figure 2: Rank collapse gives a high but trivial entropy The rank collapse of the policy network gives a policy with high entropy but zero variance across states. The network outputs the same high-entropy action distribution in all states, as all the neurons in the feature layer are dead. Its output only depends on the constant bias term.
@@
The collapse observed in the policy’s representation differs from the typical entropy collapse. As one can observe from Figure 2, the collapse in representation gives a policy that has high entropy (still lower than a uniform policy and its initialization). However, in MDPs, entropy is computed as the average across states. A high overall entropy can come from an average of high-entropy states with different action distributions or trivially from the same high-entropy distribution in all states. In our case, it is the latter: we compute the policy variance across states and observe that it collapses to zero, indicating that the policy network outputs the same action distribution at every state, which is often close to uniform. This is consistent with the policy representation, which has collapsed to a layer with mostly dead neurons  [^3] We consider a ReLU neuron as dead when its values are zero for all the samples in the batch and a tanh neuron dead when its standard deviation across samples is less than 0.001.. The output layer gives the output distribution based on only the bias term, as the linear weights are multiplied by a null feature vector. Such a policy takes the same actions across all states; therefore, performance for non-trivial tasks would collapse.
@@ 3.2 Collapsed representations create trust issues and unrecoverable loss
Intuitively, the heuristic trust region set by PPO-Clip should prevent sudden catastrophic changes and, therefore, limit the rank collapse, which induces worse performance. However, empirically, it seems the trust region is not able to mitigate the collapse. In this section, we seek to understand the interaction between the rank collapse and the trust region. We argue that as rank collapses, the clipping constraint becomes unreliable and unable to restrict learning. This is inline with previous works that have pointed out that probability ratios during training can go beyond the clipping limits with PPO-Clip (Engstrom et al., 2020; Wang et al., 2020; Sun et al., 2022). We believe, however, that this behavior is systematic when rank collapses and does not merely happen occasionally.
@@
Wang et al. (2020, Theorem 2) state that when the gradients of the unclipped samples align with the gradients of clipped samples, the clipped samples’ ratios will have their probabilities continue to go beyond the clip limit. They claim this condition would hold in practice because of “optimization tricks” or optimizer accumulated moments; however, there is no evidence that these factors induce the gradient alignment or that the alignment is present in practice. Our intuition is that representation degradation leads to alignment in the gradients and, therefore, a breakdown of the trust region constraint. This can create a snowball effect, making PPO-Clip unable to prevent representation collapse. We summarize this in two observations:
1. Loss of trust is extreme around poor representations
@@
The average of probability ratios outside the clipping limits (below $1-\epsilon$ in Figure 3) significantly diverges from the clipping limit around the collapse of the agent’s representation. This gives one more reason why the PPO trust region can be violated. We isolate this in a toy setting and analyze it formally in the next section. We further show in Figure 4 scatter plots of the lowest average probability ratios in runs with their associated representation metrics (20 points per run, across windows of size 1% training progress, spanning at least the horizon of the environment, so that points are well spaced in the run, with each point being the average of the window); we observe no significant correlation in the regions where the representation is rich (high rank, low pre-activation norm), but an apparent decrease of the average of probability ratios below $1-\epsilon$ is observed as the representation reaches poor values.
@@
Figure 3: Focusing on individual runs Individual training curves on ALE/NameThisGame-v5 with different epochs per batch. Extremely low ratios are observed around the representation collapse of a PPO-Clip agent, implying that the heuristic trust region breaks down when representation power is lacking. The last-minibatch value of the PPO objective is decreasing towards 0 around the representation collapse, implying a reduction of the ability to improve the policy and recover which is corroborated by the increase of the plasticity loss. (The ratios are trivially within the $1-\epsilon$ region after a collapse as the collapsed model does not change much, so no value below $1-\epsilon$ is logged.)
@@
Figure 4: Representation vs trust region Samples from ALE/Phoenix-v5 training curves. Each point maps an average of the probability ratios below the clipping limit vs. its corresponding average representation metric (dead neurons, feature rank, feature norm). The average ratios are significantly lower around poor representations (high dead neurons, low policy rank, high feature norm) reflecting the failure of the trust region. Averages are over non-overlapping windows larger than episodes.
2. Loss of plasticity renders performance collapse unrecoverable
@@
The monotonic decrease in performance overlaps with a monotonic decrease in policy variance and PPO objective. It appears that as the policy loses its ability to distinguish state, it can also ascend the PPO objective less and less at each batch (recall: after collecting a batch, the loss starts around zero with a normalized advantage, and through minibatch updates, the clipped policy advantage is ascended). Intuitively, this is implied by a loss of plasticity or a collapse in entropy (no new actions to learn from). As seen in Section 2 the entropy does not collapse, and measuring the plasticity loss in Figure 3 shows that that the decrease in objective gain is associated with a significant increase in plasticity loss.
@@
Hence, around collapse, the representation of the policy is getting so poor that it is impacting its ability to distinguish and act differently across states; the trust region cannot prevent this sudden and catastrophic change as it also breaks down with a poor representation; finally, the policy’s plasticity is also becoming so poor that the agent cannot recover by optimizing the surrogate objective.
@@ 3.2.1 A toy setting to understand the effects of rank collapse on trust region
Let us consider a batch containing two state-action pairs $(x,a_{1})$ and $(y,a_{1})$ with sampled probabilities $\pi_{\text{old}}(a_{1}|x)$ and $\pi_{\text{old}}(a_{1}|y)$ and positive estimated advantages $A(x,a_{1}),A(y,a_{1})>0$. Let $\phi(x),\phi(y)\in\mathbb{R}$ be fixed 1-dimensional representations of $x$, and $y$ that can be seen as the output of the (frozen) penultimate layer of a policy network with collapsed representation (all but one dead neuron),
@@
and let $\alpha\in\mathbb{R}$ such that $\phi(y)=\alpha\phi(x)$. Let ${\bm{\theta}}=[\theta_{1},\theta_{2}]$, be the last layer of the network, computing the logits of two possible actions, $a_{1}$ and $a_{2}$, that are then fed into a softmax to compute the probabilities. I.e., $\pi_{\bm{\theta}}(a_{i}|s)=\frac{e^{\theta_{1}\phi(s)}}{e^{\theta_{1}\phi(s)}+e^{\theta_{2}\phi(s)}}$. Consider PPO minibatch updates alternating between $(x,a_{1})$ and $(y,a_{1})$. Ideally, the PPO loss increases $\pi_{\bm{\theta}}(a_{1}|s)$ at gradients on $(x,a_{1})$ until it reaches the clip ratio and similarly on $(y,a_{1})$. However, we show in Appendix C that a gradient step in $(x,a_{1})$ also affects $\pi_{\bm{\theta}}(a_{1}|y)$ and depending on $\alpha$ will increase it past its clipped ratio, or decrease it below its initial value.
Essentially, when $\alpha\geq 0$, a gradient on $(x,a_{1})$ increases $\theta_{1}^{\text{new}}$ therefore increasing both $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|x)$ and $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|y)$. The same holds for a gradient on $(y,a_{1})$, causing one state to reach the clip limit first depending on $\alpha>1$ but still have the other keep pushing its probability upwards. However, when $\alpha\leq 0$, a gradient on $(x,a_{1})$ increases $\theta_{1}^{\text{new}}$ therefore increasing $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|x)$ but decreasing $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|y)$. For a gradient on $(y,a_{1})$ it is the opposite: $\theta_{1}^{\text{new}}$ decreases therefore $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|x)$ decreases and $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|y)$ increases, causing each state to reduce the probability of the other, and depending on $\alpha<1$ one of the probabilities will dominate and push the other one down. Figure 5 shows the evolution of the probabilities when simulating the updates empirically.
@@ 4 Intervening to regularize representations and non-stationarity
Having observed that PPO is affected by a consistent representation degradation that impacts its trust region heuristic and causes its performance to collapse, we turn to study interventions that aim at regularizing the representation of the policy network or reducing the non-stationarity in the optimization. We investigate whether these interventions improve the representation metrics we track and if in turn, this affects performance. We choose simple interventions that do not apply modifications to the models during training (e.g., resetting or adding neurons) or require significantly more memory (e.g., maintaining separate copies of the models). We perform interventions on the games/tasks where the collapse is the most significant. We are interested in the state of the agent at the end of the training budget. We record the performance and representation metrics for each run as averages over the last 5% of training progress. We measure the excess ratio at a timestep as the average probability ratio above $1+\epsilon$ divided by the average probability ratio below $1-\epsilon$ at that timestep. This metric gives an idea of how much the policy exceeds the trust region. Its average value is computed over the last 5% of training progress where the ratios are non-trivial, giving the same window at the end of training as the other metrics when there is no collapse, otherwise a window before total collapse covering 5% of training progress, as after collapse, the model does not change anymore and the ratios are trivially within the $1+\epsilon$ and $1-\epsilon$ limits. We give additional details on the computation of these aggregate metrics and the interventions performed in Appendix B.
PFO: Regularizing features to mitigate trust issues
@@
The motivation for our first intervention and our proposed regularization method comes from our observation that the norm of the preactivation features is consistently increasing, which can be linked to the trust issues discussed in Section 3. We seek to mitigate this effect in a way that is analogous to the PPO trust region between the optimized policy and the policy that collected the batch. We apply an $L^{2}$ loss on the difference between the pre-activated features of the optimized policy and the policy that collected the batch. This can also be thought of as a way to keep the pre-activations of the network during an update within a trust region. We apply this regularization to the pre-activations and not the activations, as dead neurons cannot propagate gradients, and even when they do, depending on the activation function, do so with a low magnitude. The regularization is an additional loss/penalty added to the overall loss. We term this loss the Proximal Feature Optimization (PFO) loss.
$L_{\pi_{\text{old}}}^{PFO}({\bm{\theta}})=\mathbb{E}_{\pi_{\text{old}}\!\!}\left[\sum_{t=0}^{t_{\max}-1}\left(\phi_{\bm{\theta}}(S_{t})-\phi_{\text{old}}(S_{t})\right)^{2}\right]$ (2)
@@
We apply two versions of PFO: one on only the penultimate layer’s pre-activations and one on all the pre-activations until the penultimate layer. In the scope of this work, we do not tune the coefficient of PFO; we pick the closest power of 10 that sets the magnitude of this loss to a similar magnitude of the clipped PPO objective tracked on the experiments without intervention. This gives a coefficient of 1 for ALE, 1 for MuJoCo with tanh, and 10 for MuJoCo with ReLU. The goal is not necessarily to obtain better performance but to see if PFO improves the representations learned by PPO and if, in turn, it affects its performance. As shown in Figure 6, the regularization of PFO effectively brings the norm of the preactivation down, the number of dead neurons down, the plasticity loss down, and the rank up. This coincides with a significant decrease in the excess probability ratio, especially in the upper tail. More importantly, we also see a significant increase in the lower tail of the returns where no collapse in performance is observed anymore on ALE/NameThisGame and ALE/Phoenix, with a slight increase in the upper tail showing that PFO can increase performance. Among the interventions we have tried, PFO provided the most consistent improvements across the tasks and metrics we monitored.
@@
Figure 6: Effects of regularizing features and non-stationarity Top: ALE/Phoenix-v5. Regularizing the difference between the features of consecutive policies with PFO results in better representations and mitigates performance collapse while improving performance in some cases. Bottom: ALE/Gravitar. Sharing the feature trunk between the actor and the critic results in a worse policy representation as the value network is subject to rank collapse due to reward sparsity. Each boxplot is formed by 15 runs including 3 different epoch configurations with 5 seeds each.
@@
In deep RL, the decision to use the same feature network trunk for both the actor and the critic is not trivial. Depending on the complexity of the environment, it can significantly vary the performance of a PPO agent (Andrychowicz et al., 2021; Huang et al., 2022a). We, therefore, attempt to draw a connection between sharing the feature trunk, the resulting representation, and its effects on the PPO objective. In this intervention, we make the actor and the critic share all the layers except their respective output layers and backpropagate the gradients from both the value and policy losses to the shared trunk. Figure 6 shows that the value loss acts as a regularizer, which decreases the feature rank, and depending on the reward’s sparsity, gives two distinct effects; In dense-reward environments such as ALE/Phoenix and ALE/NameThisGame, the ranks are concentrated at low but non-zero values: the upper tail significantly decreases compared to the baselines while the lower tail increases. This coincides with a lower feature norm, lower excess probability ratio, and, in turn, a high tail for the returns. It also increases performance in some cases. However, the opposite is true in the sparse-reward environment Gravitar: the rank completely collapses, and the feature norms and excess ratios are very high, collapsing the model’s performance. This is consistent with the observations made in the plasticity works studying value-based methods where sparse rewards deteriorate the rank of the value network. We provide training curves showing the difference in the evolution of the feature rank when sharing the actor-critic trunk in Appendix D.
Resetting the optimizer

### v3

@@ 3.1 PPO suffers from deteriorating representations
Deteriorating representation
How do the representation metrics of a PPO agent such as the feature rank and the capacity loss, evolve during training? Are they subject to the same decline observed by Kumar et al. (2021); Lyle et al. (2022) in value-based methods? Does it affect performance?
@@
As illustrated in Figure 1 with ALE/Phoenix as an example, we observe a consistent increase in the norm of the pre-activations of the feature layer of the policy network. Learning curves for all the ALE games and MuJoCo tasks considered can be found in Appendix D. The increase in feature norm is present in all the games/tasks considered in both environments, that is, with the two different model architectures and activation functions in the case of MuJoCo. We associate the rapid growth in the norm of the pre-activations with an eventual decline in the policy network’s feature rank. We observe a rank decline in five out of six ALE games and seven out of eight MuJoCo tasks (four with ReLU and three with tanh). The same observations about the increasing norm of the pre-activations can be made about the critic network. However, its rank varies more with the sparsity of the reward: in most environments, its rank experiences a significant deterioration after the policy’s performance declines (not the policy’s rank) and rewards become sparser, and in the sparse-reward game Gravitar, the critic’s rank collapses before the policy. Furthermore, capacity loss is increasing for the critic, as observed in value-based plasticity studies (Lyle et al., 2022), and we also show that is the case for the actor, for which it explodes around rank collapse.
Worse consequences
How does increasing the number of epochs per rollout to vary non-stationarity affect a PPO agent’s representation? Does it degrade as observed in DQN and SAC agents when increasing the replay ratio (Nikishin et al., 2022; Kumar et al., 2022)?
@@
Increasing the replay ratio in DQN and SAC deteriorates the agent’s representation and, in turn, its performance (Kumar et al., 2022; D’Oro et al., 2023). This is commonly attributed to “overfitting” to previous experience (Nikishin et al., 2022). Increasing the number of epochs in PPO is analogous, and a natural hypothesis is that this would accelerate the deterioration of the policy’s representation. Figure 1 shows that increasing the number of epochs accelerates the increase of pre-activations norm and the decrease of the policy’s feature rank. [^6] The collapse happens with all epochs configurations when trained for long enough as seen in Figure 32, increasing the number of epochs is a tool to observe the collapse earlier rather than a condition for it to happen. In some cases, the rank eventually collapses, coinciding with the policy’s performance collapse. We observe the performance collapse in three of the six ALE games and three of the four MuJoCo tasks.
@@
Figure 1: Deteriorating performance and representation metrics The policy network of a PPO-Clip agent on ALE/Phoenix-v5 is subject to a deteriorating representation. The norm of the pre-activations of the penultimate layer consistently increases, and its rank eventually decreases. Performing more optimization epochs per rollout to increase the effects of non-stationarity accelerates the growth of the norm of the pre-activations and the collapse of its rank. This ultimately leads to the collapse of the policy. This collapse is not driven by the value network, whose rank is still high. Both network’s ability to fit arbitrary targets (capacity loss) is also worsening.
@@
Figure 2: Rank collapse gives a high but trivial entropy The rank collapse of the policy gives a policy with high entropy but zero variance across states. The network outputs the same high-entropy action distribution in all states, as all the neurons in the feature layer are dead. Its output only depends on the constant bias term.
@@
The collapse we observe is distinct from the typical entropy collapse. Figure 2 shows that the policy reaches a high entropy. A high overall entropy can come from an average of high-entropy states with different action distributions or trivially from the same high-entropy distribution in all states. Our analysis reveals the latter, a zero policy variance across states. This corresponds to a collapsed representation where most neurons are inactive.  [^7] We consider a ReLU neuron as dead when its values are zero for all the samples in the batch and a tanh neuron dead when its standard deviation across samples is less than 0.001. The output thus relies solely on the bias term, as linear weights act on a null feature vector, making actions near uniform across all states and collapsing performance on complex tasks.
@@ 3.2 Collapsed representations create trust issues and unrecoverable loss
Intuitively, the heuristic trust region set by PPO-Clip should prevent sudden catastrophic changes and limit the rank collapse, which induces worse performance. However, empirically, it seems the trust region cannot mitigate the collapse. In this section, we seek to understand the interaction between the rank collapse and the trust region. We argue that as rank collapses, the clipping constraint becomes unreliable and unable to restrict learning. This is in line with previous works that have pointed out that probability ratios during training can go beyond the clipping limits with PPO-Clip (Engstrom et al., 2020; Wang et al., 2020; Sun et al., 2022). We believe, however, that this behavior is systematic when rank collapses and does not merely happen occasionally.
@@
Wang et al. (2020, Theorem 2) state that when the gradients of the unclipped samples align with the gradients of clipped samples, the clipped samples’ ratios will have their probabilities continue to go beyond the clip limit. They claim this condition would hold in practice because of “optimization tricks” or optimizer accumulated moments; however, there is no evidence that these factors induce the gradient alignment or that the alignment is present in practice. Our intuition is that representation degradation leads to alignment in the gradients and, therefore, a breakdown of the trust region constraint. This can create a snowball effect, preventing PPO-Clip from preventing representation collapse. We summarize this in two observations:
Loss of trust is extreme around poor representations
@@
The average of probability ratios outside the clipping limits (below $1-\epsilon$ in Figure 3) significantly diverges from the clipping limit around the collapse of the agent’s representation. This gives one more reason why the PPO trust region can be violated. We isolate this in a toy setting and analyze it formally in the next section. We further show in Figure 4 scatter plots of the lowest average probability ratios in runs with their associated representation metrics.88820 points per run, across windows of size 1% training progress, spanning at least the horizon of the environment, so that points are well spaced in the run, with each point being the average of the window We observe no significant correlation in the regions where the representation is rich (high rank, low pre-activation norm), but an apparent decrease of the average of probability ratios below $1-\epsilon$ is observed as the representation reaches poor values. Note that we characterize the collapsing regime by an extremely low rank, however, it is not straightforward to draw a line between low-rank representations beneficial for generalization and extremely low-rank representations causing aliasing as also acknowledged by Gulcehre et al. (2022), but for environments like Atari, our figures seem to draw the line at single-digit ranks, which can be related to the action space of dimension 8+.
@@
Figure 3: Focusing on individual runs Individual training curves on ALE/NameThisGame-v5 with different epochs per batch. Extremely low ratios are observed around the representation collapse of a PPO-Clip agent, implying that the heuristic trust region breaks down when representation power is lacking. The last-minibatch value of the PPO objective decreases towards 0 around the representation collapse, implying a reduction in the ability to improve the policy and recover, which is corroborated by the increase in capacity loss. (Ratios are trivially above $1-\epsilon$ after collapse as a collapsed model does not change much to have values below $1-\epsilon$.)
@@
Figure 4: Representation vs trust region Samples from ALE/Phoenix-v5 training curves. Each point maps an average of the probability ratios below the clipping limit vs. its corresponding average representation metric (dead neurons, feature rank, feature norm). The average ratios are significantly lower around poor representations (high dead neurons, low policy rank, high feature norm) reflecting the failure of the trust region in this regime. Averages are over non-overlapping windows larger than episodes.
Loss of plasticity makes performance collapse unrecoverable
@@
The persistent decrease in performance overlaps with a monotonic decrease in policy variance and PPO objective. It appears that as the policy loses its ability to distinguish state, it can also ascend the PPO objective less and less at each batch (recall: after collecting a batch, the loss starts around zero with a normalized advantage, and through minibatch updates, the clipped policy advantage is ascended). Intuitively, this is implied by a loss of plasticity or a collapse in entropy (no new actions to learn from). As seen in Section 2 the entropy does not collapse, and measuring the capacity loss in Figure 3 shows that the decrease in objective gain is associated with a significant increase in capacity loss, implying loss of plasticity.
@@
Hence, around collapse, the representation of the policy is getting so poor that it is impacting its ability to distinguish and act differently across states; the trust region cannot prevent this catastrophic change as it also breaks down with a poor representation; finally, the policy’s plasticity is also becoming so poor that the agent cannot recover by optimizing the surrogate objective.
Implications and discussion
The causal connection we draw between the representation dynamics, the trust region, and the performance primarily holds around the collapse regime and not necessarily throughout training. However, this does not mean that one should only be concerned about the link when performance is starting to deteriorate. The representations don’t collapse all of a sudden; they deteriorate throughout training until they reach collapse. Thus, mitigating representation degradation should happen throughout training and not only when around the collapsing regime. In addition, the connection gives important insights into the failure mode of the popular PPO-Clip algorithm, whose trust region is highly dependent on the representation quality, and more generally about trust-region methods which only constrain the output probabilities.
@@ 3.2.1 A toy setting to understand the effects of rank collapse on trust region
Let us consider a batch containing two state-action pairs $(x,a_{1})$ and $(y,a_{1})$ with sampled probabilities $\pi_{\text{old}}(a_{1}|x)$ and $\pi_{\text{old}}(a_{1}|y)$ and positive estimated advantages $A(x,a_{1}),A(y,a_{1})>0$. Let $\phi(x),\phi(y)\in\mathbb{R}$ be fixed 1-dimensional representations of $x$, and $y$ that can be seen as the output of the (frozen) penultimate layer of a policy network with collapsed representation (all but one dead neuron), and let $\alpha\in\mathbb{R}$ such that $\phi(y)=\alpha\phi(x)$. Let ${\bm{\theta}}=[\theta_{1},\theta_{2}]$, be the last layer of the network, computing the logits of two actions, $a_{1}$ and $a_{2}$, that are fed into a softmax to compute the probabilities. I.e., $\pi_{\bm{\theta}}(a_{i}|s)=\frac{e^{\theta_{i}\phi(s)}}{e^{\theta_{1}\phi(s)}+e^{\theta_{2}\phi(s)}}$. Consider PPO minibatch updates alternating between $(x,a_{1})$ and $(y,a_{1})$. Ideally, the PPO loss increases $\pi_{\bm{\theta}}(a_{1}|s)$ at gradients on $(x,a_{1})$ until it reaches the clip ratio and similarly on $(y,a_{1})$. However, we show in Appendix C that a gradient step in $(x,a_{1})$ also affects $\pi_{\bm{\theta}}(a_{1}|y)$ and depending on $\alpha$ will increase it past its clipped ratio, or decrease it below its initial value. Essentially, when $\alpha\geq 0$, a gradient on $(x,a_{1})$ increases $\theta_{1}^{\text{new}}$ therefore increasing both $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|x)$ and $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|y)$. The same holds for a gradient on $(y,a_{1})$, causing one state to reach the clip limit first depending on $\alpha>1$ but still have the other keep pushing its probability upwards. However, when $\alpha\leq 0$, a gradient on $(x,a_{1})$ increases $\theta_{1}^{\text{new}}$ therefore increasing $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|x)$ but decreasing $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|y)$. For a gradient on $(y,a_{1})$ it is the opposite: $\theta_{1}^{\text{new}}$ decreases therefore $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|x)$ decreases and $\pi_{{\bm{\theta}}^{\text{new}}}(a_{1}|y)$ increases, causing each state to reduce the probability of the other, and depending on $\alpha<1$ one of the probabilities will dominate and push the other one down. Figure 5 shows the evolution of the probabilities when simulating the updates empirically.
@@ 4 Intervening to regularize representations and non-stationarity
Having observed that PPO is affected by a frequent representation degradation that impacts its trust region heuristic and causes its performance to collapse, we turn to study interventions that aim at regularizing the representation of the policy network or reducing the non-stationarity in the optimization. We investigate whether these interventions improve the representation metrics we track and if in turn, this affects performance. We choose simple interventions that do not apply modifications to the models during training (e.g., resetting or adding neurons) or require significantly more memory (e.g., maintaining separate copies of the models). We perform interventions on the games/tasks where the collapse is the most significant. We are interested in the state of the agent at the end of the training budget. We record the performance and representation metrics for each run as averages over the last 5% of training progress. We measure the excess ratio at a timestep as the average probability ratio above $1+\epsilon$ divided by the average probability ratio below $1-\epsilon$ at that timestep. This metric gives an idea of how much the policy exceeds the trust region. Its average value is computed over the last 5% of training progress where the ratios are non-trivial, giving the same window at the end of training as the other metrics when there is no collapse, otherwise a window before total collapse covering 5% of training progress, as after collapse, the model does not change anymore and the ratios are trivially within the $1+\epsilon$ and $1-\epsilon$ limits. We give additional details on the computation of these aggregate metrics and the interventions performed in Appendix B.
PFO: Regularizing features to mitigate trust-region issues
@@
The motivation for our first intervention and our proposed regularization method comes from our observation that the norm of the preactivation features is consistently increasing, which can be linked to the trust-region issues discussed in Section 3. We seek to mitigate this effect in a way that is analogous to the PPO trust region, by extending the trust region to the feature space. We apply an $L^{2}$ loss on the difference between the pre-activated features of the optimized policy and the policy that collected the batch, as a way to keep the pre-activations of the network during an update within a trust region. We apply this regularization to the pre-activations and not the activations, as dead neurons cannot propagate gradients, and even when they do, depending on the activation function, do so with a low magnitude. The regularization is an additional loss/penalty added to the overall loss. We term this loss the Proximal Feature Optimization (PFO) loss. With $\phi_{\bm{\theta}}(s)$ as the pre-activation of the penultimate layer of the actor $\pi_{\bm{\theta}}$ given a state $s$,
$L_{\pi_{\text{old}}}^{PFO}({\bm{\theta}})=\mathbb{E}_{\pi_{\text{old}}\!\!}\left[\sum_{t=0}^{t_{\max}-1}\lVert\phi_{\bm{\theta}}(S_{t})-\phi_{\pi_{\text{old}}}(S_{t})\rVert^{2}_{2}\right].$ (2)
@@
We apply two versions of PFO: one on only the penultimate layer’s pre-activations and one on all the pre-activations until the penultimate layer. In the scope of this work, we do not tune the coefficient of PFO; we pick the closest power of 10 that sets the magnitude of this loss to a similar magnitude of the clipped PPO objective tracked on the experiments without intervention. This gives a coefficient of 1 for ALE, 1 for MuJoCo with tanh, and 10 with ReLU. The goal is not necessarily to obtain better performance but to see if PFO improves the representations learned by PPO and if, in turn, it affects its trust region and performance. As shown in Figure 6, the regularization of PFO effectively brings the norm of the preactivation down, the number of dead neurons down, the capacity loss down, and the rank up. This coincides with a significant decrease in the excess probability ratio, especially in the upper tail. More importantly, we also see a significant increase in the lower tail of the returns where no collapse in performance is observed anymore on ALE/NameThisGame and ALE/Phoenix, with a slight increase in the upper tail showing that PFO can increase performance. Among the interventions we have tried, PFO provided the most consistent improvements in representation and trust region.
@@
Figure 6: Effects of regularizing features and non-stationarity Top & Middle: ALE/Phoenix-v5 & ALE/NameThisGame-v5. Regularizing the difference between the features of consecutive policies with PFO results in better representations, a lower trust-region excess, and mitigates performance collapse. The same applies to sharing the actor-critic trunk. Bottom: ALE/Gravitar. Sharing the feature trunk between the actor and the critic results in a worse policy representation as the value network is subject to rank collapse due to reward sparsity. A boxplot includes 15 runs with different epochs.
@@
In deep RL, the decision to use the same feature network trunk for both the actor and the critic is not trivial. Depending on the complexity of the environment, it can significantly change the performance of a PPO agent (Andrychowicz et al., 2021; Huang et al., 2022a). We, therefore, attempt to draw a connection between sharing the feature trunk, the resulting representation, and its effects on the PPO objective. In this intervention, we make the actor and the critic share all the layers except their respective output layers and backpropagate the gradients from both the value and policy losses to the shared trunk. Figure 6 shows that the value loss acts as a regularizer, which decreases the feature rank and, depending on the reward’s sparsity, gives two distinct effects. In dense-reward environments such as ALE/Phoenix and ALE/NameThisGame, the ranks are concentrated at low but non-trivial values: the upper tail significantly decreases compared to the baselines while the lower tail increases. This coincides with a lower feature norm, lower excess probability ratio, and, in turn, a high tail for the returns. It also increases performance in some cases. However, the opposite is true in the sparse-reward environment Gravitar: the rank completely collapses, and the feature norms and excess ratios are very high, collapsing the model’s performance. This is consistent with the observations made in the plasticity works studying value-based methods: they show that sparse rewards deteriorate the rank of the value network, and we show that when shared in an actor-critic architecture they, in turn, deteriorate the policy. It is important to note that this distinction using the reward sparsity holds when comparing environments from the same family (e.g., ALE), but may not hold otherwise (e.g., comparing an ALE and a MuJoCo environment). We provide training curves showing the difference in the evolution of the feature rank when sharing the actor-critic trunk in Appendix D. To further strengthen this observation we run an intervention on ALE/Phoenix (a dense reward environment), with a reward mask randomly masking a reward with 90% chance, comparing the effects of sharing the actor-critic trunk. As expected, while with dense rewards, sharing the trunk is beneficial in ALE/Phoenix (Appendix Figure 21), with the sparse reward, the opposite is true: sharing the trunk is detrimental (Appendix Figure 35).
Adapting Adam